OpenStudy (anonymous):

(x-4) - (x + 7) + (x+3) 7y 8y 6y * problem read like so: x - 4 over 7y minus x + 7 over 8y plus x+ 3 over 6y this problem has 7 empty places for the numerator and 4 places for the denominator.. . . every answer i get does not fit in the required space. . . please explain.

OpenStudy (mimi_x3):

is it: \[\large \frac{x-4}{7y} -\frac{x+7}{8y} +\frac{x+3}{6y} \]?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

Whatdo you want to do with it?

OpenStudy (anonymous):

i have to simplify it but i cant get the right answer

OpenStudy (anonymous):

Is\[\frac{31 x-159}{168 y} \]the right answer?

OpenStudy (anonymous):

idk if it is bc the automatic checker isnt working can u explain?

OpenStudy (anonymous):

Combine the first two fractions. Then combine the result with the third fraction.\[\frac{x-4}{7y }-\frac{x+7}{8y}+\frac{x+3}{6y} \]\[\frac{x-81}{56 y}+\frac{x+3}{6 y} \]\[\frac{31 x-159}{168 y} \]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

You are welcome.